a gaseous organic compound has a density of 2.5 kg per metre cube at 2 ATM and at 273 degree Celsius the molecular formula of the compound can be

Grade:11

1 Answers

Arun
5 years ago
PV=nRT
 
D=m/V
Molar Mass=m/n
so
m=DV
m=nMM
so
DV=nMM
so
V=nMM/D
 
Plug into PV=nRT
 
PnMM/D=nRT
n on both sides cancel
 
PMM/D=RT
 
Solve for MM
 
MM=DRT/P
 
Plug in
 
hope it helps
 
Regards
Arun (askIITians Forum expert)